About 5​% of hourly paid workers in a region earn the prevailing minimum wage or less. A grocery chain offers discount rates to companies that have at least 30 employees who earn the prevailing minimum wage or less.Company B has 542 employees. What is the probability that Company B will get the​ discount?
Answer #1

Let X is a radnom variable shows the number of employees out of 542 earn the prevailing minimum wage or less.Here X has binomial distribution with parameters n=542 and p=0.05.

Since np and n(1-p) are greater than 5 so using normal approximation X has normal distribution with parameters

u= n*p=542 * 0.05 = 27.1

and standard deviation

0= Vn*p*(1-p) = 5.0739

here we need to use continuity correction. The z-score for X = 30-0.5 = 29.5 is

| 29.5-27.1 5,0379 - = 0.4730

The probability that Company B will get the​ discount is

P(X > 30) = P(Z > 0.4730) = 0.3181

The excel function used for finding area right to z-score 0.4730 is " =1-NORMSDIST(0.4730".

Answer: 0.3181